1. Suppose you are the manager of a Golf Club with monopoly power. A typical consumer’s

inverse demand function for your firm’s product is 𝑃 = 100 − 20𝑄, and your cost function is C(Q) = 20Q.

(a) If you apply a two-part pricing strategy, how much would be the membership fee and the per unit price?

(b) How much additional profit do you earn using a two-part pricing strategy compared with single price strategy?


1
Expert's answer
2022-03-01T10:52:36-0500

"\ud835\udc43 = 100 \u2212 20\ud835\udc44"

"C(Q) = 20Q"

a) Applying two pricing strategies the per-unit price becomes the price at which price equals the marginal cost of production.

P=MC

100−20Q=20=4

P= 20

Membership fees= the area of the triangle above the MC curve, and the per-unit price becomes the price

"\u0394\n=\n1\n2\n(\n100\n\u2212\n20\n)\n(\n4\n\u2212\n0\n)\n\u0394\n=\n160"

b. If there was a single pricing strategy, then equilibrium will occur at MR = MC

MR=MC

100−40Q=20Q=2

P=60

The profit would have been

π"_1" =TR−TC

π"_1" =60×2−20×2

π"_1" =120−40

π"_1" =80

When there is two pricing strategy, profit is given as:

π"_2" =(20×4+160)−20×4

π"_2"=80+160−80

π"_2"=160

Additional profit will be:

Δπ=π"_2"−π"_1" =160−80=80